# Reality’s Canvas, Language’s Brush: Crafting 3D Avatars from Monocular Video

### 08 Dec 2023 v1

@@ Abstract
Recent advancements in 3D avatar generation excel with multi-view supervision for photorealistic models. However, monocular counterparts lag in quality despite broader applicability. We propose ReCaLaB to close this gap. ReCaLaB is a fully-differentiable pipeline that learns high-fidelity 3D human avatars from just a single RGB video. A pose-conditioned deformable NeRF is optimized to volumetrically represent a human subject in canonical T-pose. The canonical representation is then leveraged to efficiently associate viewpoint-agnostic textures using 2D-3D correspondences. This enables to separately generate albedo and shading which jointly compose an RGB prediction. The design allows to control intermediate results for human pose, body shape, texture, and lighting with text prompts. An image-conditioned diffusion model thereby helps to animate appearance and pose of the 3D avatar to create video sequences with previously unseen human motion. Extensive experiments show that ReCaLaB outperforms previous monocular approaches in terms of image quality for image synthesis tasks. ReCaLaB even outperforms multi-view methods that leverage up to 19x more synchronized videos for the task of novel pose rendering. Moreover, natural language offers an intuitive user interface for creative manipulation of 3D human avatars.
@@ 1 Introduction
Enter ReCaLaB, a gateway where monocular 3D avatar creation converges with an intuitive language interface. Leveraging a single monocular video, we produce a digital 3D human that can be effortlessly manipulated with a user text prompt. This liberates avatar creation from the confines of specialized lab setups, eliminating the need for hardware expertise to construct complex camera arrays while expert coding knowledge is no longer required to manipulate learned digital representations. The title "Reality’s Canvas, Language’s Brush" encapsulates this vision, where language acts as the brushstroke that brings avatars to life on the canvas of real monocular video data, fostering creative expression on a global scale and empowering a broader community.
Technically, we leverage the recent progress in implicit 3D neural radiance fields [34] and their deformable counterparts targeting 3D human generation [20, 60, 8] to generate a 3D avatar. Our pipeline ReCaLaB is illustrated in Fig. 1. At its core, ReCaLaB employs a fully-differentiable pipeline. Leveraging the seminal ideas of HumanNeRF [60], we learn a deformation field from a monocular video that transports pixel input into a canonical 3D representation conditioned on the human skeletal motion in the video frames. Using learned 3D-2D correspondences from this canonical T-pose, we generate viewpoint-agnostic neural texture stacks [8] representing separately albedo RGB and shading to efficiently compose a final 2D rendering. A textual prompt can then be leveraged by a diffusion model [57, 5] to generate either a human motion or change geometry and appearance of the 3D avatar. The fully-differentiable nature of this pipeline thereby allows to propagate gradients of an image-conditioned diffusion model [5] for a targeted change of individual components in ReCaLaB. Controlled un-freezing of intermediate parts enables change of body shape, texture, or lighting while being free to choose the camera viewpoint.
The result is a model that not only outperforms existing monocular approaches in terms of image quality but also surpasses multi-view methods, even when the latter leverage a significantly greater volume of data.
@@
    A Viewpoint-agnostic Canonical Representation of 2D textures and 2D-3D correspondences allows to learn high-fidelity 3D textures from a single RGB video source. Resulting rendering quality under novel viewpoints and human poses outperforms even multi-view approaches.
@@
    A Decomposed Appearance Generation module discerns albedo and shading predictions. This enables modification of illumination while preserving texture integrity.
Figure 1: Framework of ReCaLaB. Our approach takes monocular video frames as the input to reconstruct the human avatar. We first learn the backward deformation $T$ of the human body in module (a) with the human pose. Then in module (b), we generate a corresponding UVS map by using a volume generator $G_{c}^{F}$ initialized by the SMPL T-pose prior, followed with different mapping functions $G_{c}^{M},(M\in\{U,V,S\})$. We then designed a texture generator $H_{c}^{T}$ along with the albedo branch $H_{c}^{A}$, and shading branch $H_{c}^{N}$ for the neural texture learning in module (c). Finally, in module (d), shape, texture, and shading modules can be updated accordingly with the instruction from the language brush.
@@ 2 Related Works
In the realm of 3D human representation, progress has been notable, with early methods focusing on predicting keypoints and skeletons [6, 16], later evolving to incorporate volumetric details from data [32] including details for hands, face, and feet [43, 38]. Dataset acquisition has seen substantial efforts since early multi-view reconstruction works [22], with advancements in hardware setups and accurate annotations [17, 45, 65, 18]. However, obtaining high-quality 3D annotations remains a costly endeavor [58, 21, 18]. In the pursuit of animatable human avatars, the field has progressed from static 3D surface recovery [48, 49] to sophisticated neural approaches guided by 3D human models. The increasing popularity of implicit scene representations [33, 39], in particular static neural fields [34, 35, 3] and their deformable counterparts [29, 46] have accelerated this process even from monocular video input [40, 41, 30]. The emergence of neural rendering methods, including Neural Body [46], Neural Articulated Radiance Field [37], A-NeRF [53], TAVA [27], and PoseVocab [31], signify a significant shift towards learning human shape, appearance, and pose from multi-view data. These methods allow to synthesize photorealistic images from novel 3D viewpoints and to animate human characters with given poses. Additionally, for monocular video, approaches like SelfRecon [19] and NeuMan [20] as well as Animatable NeRF [44] and HumanNeRF [60] integrate motion priors for regularization to allow this even with a single input video. While these innovations excel in rendering quality, they come with the challenge of high computational demands during both training and testing phases. Additionally, they lack the capability for intuitive texture editing due to their implicit scene representation. Our pipeline also generate a detailed human avatar from just a single monocular video, but additionally offers an intuitive way to manipulate its texture, pose, and shape through text prompts. This unique capability is made possible by leveraging 2D-3D correspondences.
@@
NeRFs excel in reproducing visual appearance even with fixed geometry [7]. While previous approaches improve texture fidelity through adversarial training [59] or latent diffusion [56], correspondences maps to 3D human models in the form of SMPL [32] or DensePose [11] are a nifty alternative. UV Volumes [8] leverages this strength to create an efficient rendering pipeline for human avatars, utilizing 2D-3D correspondences, a technique also employed in other domains [26]. This significantly accelerates both learning and inference processes while allowing to directly manipulate texture in 2D. The method shows impressive results trained when being trained with multi-view data. We also leverage a correspondence map in the form of 24 human body segments, and learn a canonical UV-map together with a neural texture stack to capitalize on information from just a monocular video mapped into a canonical space.
@@ 2.2 Manipulation with Diffusion Models
Diffusion models, a class of generative models employing stochastic diffusion [50, 52], have shown promise in various applications. During training, noise is gradually introduced to a sample from the data distribution, while the inverse process is learned starting from pure initial noise. This technique has been extended for image generation [14, 51], and in recent years, conditioned diffusion models have gained traction in text-guided image generation [9, 36, 13], where an image guides the denoising process. In the realm of human motion, diffusion models have been leveraged to generate complex motion sequences for human avatars [62]. We adopt a similar diffusion model [57] to generate text-driven motion sequences within ReCaLaB.
Noteworthy examples in human image generation include DreamPose [23], which creates synthetic fashion videos guided by still images and pose sequences, and ControlNet [64], capable of generating human images based on novel poses and textual descriptions. Works like Dreamface [63] and TeCH [15] employ diffusion models to learn editable textures for synthetic human faces and bodies. Moreover, InstructPix2Pix [5] stands out for its impressive results, and its extension, Instruct-NeRF2NeRF [12], ensures frame consistency in static real-world scenes. In our approach, we apply a similar diffusion process to a deformable scene mapped to a canonical parametric human representation, allowing us to edit a generated stack of 2D albedo and shading maps conditioned on a text prompt.
@@ 3 Method
Given a monocular video that captures the human body, we propose an efficient technique to reconstruct the 3D human avatar, which can be further driven by a single-sentence instruction. The overview of our method is shown in Fig. 1, where our approach mainly consists of three fundamental components: 1) we learn a hybrid neural representation to model the human avatar, with a neural feature volume field and neural texture extraction. 2) We introduce a shading network to refine the learned neural texture to cope with lighting differences. 3) In the end, we update the editable neural human representation with language instruction as a brush. With our full pipeline, we can lift the human into a 3D editable neural avatar from a monocular human video alone and edit it with human language descriptions.
@@ 3.1 Reality’s Canvas: Dynamic Human Modeling
$F_{c}:(\mathbf{x}_{c},\mathbf{d})\mapsto\mathbf{c},$ (1)
where $\mathbf{x}_{c}$ is a sampled 3D point from the human body in canonical space and another input, $\mathbf{d}$, is the viewing direction. $\mathbf{c}$ is the mapped color in texture space output by the mapping function, $F_{c}$.
@@
where $G_{c}^{F}$ servers as the mapping function to take the 3D input position $\mathbf{x}_{c}$ to obtain the $\mathbf{f}$ to form the neural feature volume.
@@
$G_{c}^{U}:(\mathbf{f},\mathbf{d})\mapsto\mathbf{u},\;\;G_{c}^{V}:(\mathbf{f},\mathbf{d})\mapsto\mathbf{v}$ (3)
@@
$G_{c}^{S}:(\mathbf{f},\mathbf{d})\mapsto\mathbf{s}$ (4)
To model the dynamics in human motion and animate the lifted 3D human avatar, we further introduce the deformation module, $T$, to connect the 3D points $\mathbf{x}$ in the observation space, with the one in the canonical space, $\mathbf{x}_{c}$. In other words, the mapping function in the observation space is connected to the aforementioned mapping function in canonical space via $T$. Thus, we can obtain the neural volume in observation space, by warping the canonical neural volume:
$G_{c}^{S}(\mathbf{f},\mathbf{d})=G_{c}^{S}(G_{c}^{f}(T(\mathbf{x},\mathbf{P})),\mathbf{d}),$ (5)
where $\mathbf{P}$ denotes the human pose for the given time frame. We apply the same deformation field to bend the rays of the mapping function for UV representations. Volume rendering is then applied to the output $\mathbf{U}$, $\mathbf{V}$ and the semantic map $\mathbf{S}$.
@@ 3.2 Disentangled Neural Texture Learning
With the rendered UV map and semantic map, our network can further synthesize high-quality neural textures. We realize this by decomposing appearance generation into albedo and shading components within a disentangled neural texture learning module. Specifically, we build another neural mapping function, $H_{c}^{T}$, to generate the neural texture field in the texture space with regard to the input UV and semantic space:
@@
Compared to learning a neural radiance field to synthesize the color of the human avatar, learning a neural texture will produce better appearance modeling by avoiding all the floaters above the texture surfaces. However, neural texture cannot reflect the view-dependent color effect when we change the viewpoint, nor the pose-dependent effect. To cope with this issue, we then decompose the neural texture representation into the albedo component and shading information. Firstly, again built by MLP, takes the neural texture representation and converts to the albedo color component, reflecting the diffused color of the surface:
@@
where $\mathbf{a}$, denotes the albedo color componet and, $H_{c}^{A}$, represents the mapping function in the albodo branch.
We further encode neural shading information to represent the lighting condition influenced by the pose articulation and view direction change:
@@
where $H_{c}^{N}$ is the learned neural mapping function, that maps the neural texture, $\mathbf{t}$, with the guidance of view direction, $\mathbf{d}$ and human poses, $\mathbf{P}$, to the shading information, $\mathbf{m}$. To this end, the final color for each ray in 2D image plane is calculated as $\mathbf{c}=\mathbf{a}\cdot\mathbf{m}$.
@@ 3.3 Languages’s Brush: Editing Neural Human
Our disentangled neural texture generation allows us to manipulate texture and shape separately. This gives us explicit control over which part of the pipeline is changed for synthesis upon user inputs. Given text instruction for neural human editing, our framework can be adapted to update the corresponding module related to the instruction, enabling more accurate and efficient neural human editing.
@@
Figure 3: Qualitative results on novel poses guided by text instruction. a) and b) are frames from the instructions "A person turns to his right and paces back and forth", and "A person is running" respectively. ReCaLaB provides sharper texture and fewer geometric artifacts compared to HumanNeRF [60].
@@ 3.4 Training Objectives
where $\lambda_{\text{LPIPS}}$ and $\lambda_{\text{MSE}}$ are the weights for the corresponding losses. The photometric loss will also back-propagate to train the mapping functions, $G_{c}^{U}$, $G_{c}^{V}$, and $G_{c}^{S}$ to predict meaningful UV and semantic maps, which are further regularised by another three losses, $\mathcal{L}_{reg}^{U}$, $\mathcal{L}_{reg}^{V}$ and $\mathcal{L}_{reg}^{S}$ between the predictions and pseudo labels, obtained from a pre-trained Densepose [11] network. In addition to the regularization losses, we add a smoothnes term $\mathcal{L}_{\text{smt}}$ to encourage the neighbor 3D points to embed the same semantics by leveraging human parsing priors.
@@
where both $\mathcal{L}_{\text{reg}}^{U}$ and $\mathcal{L}_{\text{reg}}^{V}$ are formulated with MSE loss, while $\mathcal{L}_{\text{reg}}^{S}$ uses cross entropy loss. To reduce flying pixels and condense the learned neural volume in line with the human shape, we utilize a mask loss to constrain geometry reconstruction with $\mathcal{L}_{\text{mask}}$, similar to [61]. Finally, we obtain the overall training objective in Eq. 11 by weighting all the aforementioned losses to generate the neural human avatar:
@@
where $\lambda_{\text{rec}}$, $\lambda_{\text{reg}}$ and $\lambda_{\text{mask}}$ are the loss weight for $\mathcal{L}_{\text{rec}}$, $\mathcal{L}_{\text{reg}}$ and $\mathcal{L}_{\text{mask}}$, respectively. Also note that $\mathcal{L}_{\text{rec}}$ is used when language instruction is available for editing, while the edited 2D images will be gradually involved in the supervision.

### 24 Mar 2024 v2

@@ Abstract
Recent advancements in 3D avatar generation excel with multi-view supervision for photorealistic models. However, monocular counterparts lag in quality despite broader applicability. We propose ReCaLaB to close this gap. ReCaLaB is a fully-differentiable pipeline that learns high-fidelity 3D human avatars from just a single RGB video. A pose-conditioned deformable NeRF is optimized to volumetrically represent a human subject in canonical T-pose. The canonical representation is then leveraged to efficiently associate neural textures using 2D-3D correspondences. This enables the separation of diffused color generation and lighting correction branches that jointly compose an RGB prediction. The design allows to control intermediate results for human pose, body shape, texture, and lighting with text prompts. An image-conditioned diffusion model thereby helps to animate appearance and pose of the 3D avatar to create video sequences with previously unseen human motion. Extensive experiments show that ReCaLaB outperforms previous monocular approaches in terms of image quality for image synthesis tasks. Moreover, natural language offers an intuitive user interface for creative manipulation of 3D human avatars.
Keywords:
Neural Radiance Field Diffusion Model 3D digital human
@@ 1 Introduction
Figure 1: Overview of ReCaLaB. We propose ReCaLaB, a novel approach for photo-realistic 3D avatar creation from monocular video and different types of manipulations. Our approach models neural humans with a disentangled neural texture field and enables different types of manipulation, including novel view rendering, novel pose animation, and novel appearance with customized shape, texture, and illumination by language brush.
@@
Enter ReCaLaB, a gateway where monocular 3D avatar creation converges with an intuitive language interface. Leveraging a single monocular video, we produce a digital 3D human that can be effortlessly manipulated with a user text prompt. This liberates avatar creation from the confines of specialized lab setups, eliminating the need for hardware expertise to construct complex camera arrays while expert coding knowledge is no longer required to manipulate learned digital representations. The title “Reality’s Canvas, Language’s Brush” encapsulates this vision, where language acts as the brushstroke that brings avatars to life on the canvas of real monocular video data, fostering creative expression on a global scale and empowering a broader community.
Technically, we leverage the recent progress in implicit 3D neural radiance fields [41] and their deformable counterparts targeting 3D human generation [22, 70, 9] to generate a 3D avatar. Our pipeline ReCaLaB is illustrated in Fig. 2. At its core, ReCaLaB employs a fully-differentiable pipeline. Leveraging the seminal ideas of HumanNeRF [70], we learn a deformation field from a monocular video that transports pixel input into a canonical 3D representation conditioned on the human skeletal motion in the video frames. Using learned 3D-2D correspondences from this canonical T-pose, we generate viewpoint-agnostic neural texture stacks [9] to represent the diffused color, along with the lighting correction to efficiently compose a final 2D rendering. A textual prompt can then be leveraged by a diffusion model [66, 6] to generate either a human motion or change the geometry and appearance of the 3D avatar. The fully-differentiable nature of this pipeline thereby allows to propagate gradients of an image-conditioned diffusion model [6] for a targeted change of individual components in ReCaLaB. Controlled un-freezing of intermediate parts enables changing body shape, texture, or lighting while being free to choose the camera viewpoint.
The result is a model that not only outperforms existing monocular approaches in terms of image quality but also surpasses multi-view methods, even when the latter leverages a significantly greater volume of data.
@@
    A Viewpoint-agnostic Canonical Representation of 2D diffused textures and 2D-3D correspondences allows to learn high-fidelity 3D textures from a single RGB video source. Resulting in rendering quality under novel viewpoints and human poses outperforms even multi-view approaches.
@@
    A Decomposed Appearance Generation module discerns diffused color and lighting correction branches based on the pose and view-direction predictions. This enables modification of illumination while preserving texture integrity.
@@ 2 Related Works
In the realm of 3D human representation, progress has been notable, with early methods focusing on predicting keypoints and skeletons [7, 18], later evolving to incorporate volumetric details from data [38] including details for hands, face, and feet [50, 46]. Dataset acquisition has seen substantial efforts since early multi-view reconstruction works [25], with advancements in hardware setups and accurate annotations [19, 53, 76, 20]. However, obtaining high-quality 3D annotations remains a costly endeavor [67, 24, 20]. In the pursuit of animatable human avatars, the field has progressed from static 3D surface recovery [57, 58] to sophisticated neural approaches guided by 3D human models. The increasing popularity of implicit scene representations [39, 47], in particular static neural fields [41, 43, 4] and their deformable counterparts [35, 54] have accelerated this process even from monocular video input [48, 49, 37]. The emergence of neural rendering methods, including Neural Body [54], Neural Articulated Radiance Field [45], A-NeRF [62], TAVA [33], and PoseVocab [36], signify a significant shift towards learning human shape, appearance, and pose from multi-view data. These methods allow to synthesize photorealistic images from novel 3D viewpoints and to animate human characters with given poses. Additionally, for monocular video, approaches like SelfRecon [21] and NeuMan [22] as well as Animatable NeRF [52] and HumanNeRF [70] integrate motion priors for regularization to allow this even with a single input video. While these innovations excel in rendering quality, they come with the challenge of high computational demands during both training and testing phases. Concurrent to our approach, explicit methods using deformable 3D Gaussian splatting [27] is also explored for 3D avatar generation [23, 42, 30]. Both directions, however, lack the capability for intuitive texture editing due to their 3D scene representation. Our pipeline also generates a detailed human avatar from just a single monocular video, but additionally offers an intuitive way to manipulate its texture, pose, and shape through text prompts. This unique capability is made possible by leveraging 2D-3D correspondences.
@@
NeRFs excel in reproducing visual appearance even with fixed geometry [8]. While previous approaches improve texture fidelity through adversarial training [69] or latent diffusion [65], correspondences maps to 3D human models in the form of SMPL [38] or DensePose [12] are a nifty alternative. UV Volumes [9] leverages this strength to create an efficient rendering pipeline for human avatars, utilizing 2D-3D correspondences, a technique also employed in other domains [32]. This significantly accelerates both learning and inference processes while allowing to directly manipulate texture in 2D. The method shows impressive results when being trained with multi-view data. We also leverage a correspondence map in the form of 24 human body segments, and learn a canonical UV-map together with a neural texture stack to capitalize on information from just a monocular video mapped into a canonical space.
@@ 2.2 Manipulation with Diffusion Models
Diffusion models, a class of generative models employing stochastic diffusion [59, 61], have shown promise in various applications. During training, noise is gradually introduced to a sample from the data distribution, while the inverse process is learned starting from pure initial noise. This technique has been extended for image generation [14, 60], mesh editing [40, 55], and in recent years, conditioned diffusion models have gained traction in text-guided image generation [10, 44, 15], where an image guides the denoising process. In the realm of human motion, diffusion models have been leveraged to generate complex motion sequences for human avatars [73]. We adopt a similar diffusion model [66] to generate text-driven motion sequences within ReCaLaB.
Noteworthy examples in human image generation include DreamPose [26], which creates synthetic fashion videos guided by still images and pose sequences, and ControlNet [75], capable of generating human images based on novel poses and textual descriptions. Works like Dreamface [74], TeCH [17], and [31, 3, 16, 68, 71] employ diffusion models to learn editable textures for synthetic human faces and bodies. Moreover, InstructPix2Pix [6] stands out for its impressive results, and its extension, Instruct-NeRF2NeRF [13], ensures frame consistency in static real-world scenes. In our approach, we apply a similar diffusion process to a deformable scene mapped to a canonical parametric human representation, allowing us to edit the UVS map or neural texture stack conditioned on a text prompt.
Figure 2: Framework of ReCaLaB. Our approach takes monocular video frames as the input to reconstruct the human avatar. We first learn the backward deformation $T$ of the human body in module (a) with the human pose. Then in module (b), we generate a corresponding UVS map by using a volume generator $G_{c}^{F}$ initialized by the SMPL T-pose prior, followed with different mapping functions $G_{c}^{M},(M\in\{U,V,S\})$. We then designed a texture generator $H_{c}^{T}$ along with the diffused color branch $H_{c}^{A}$, and the lighting correction branch $H_{c}^{N}$ for the neural texture learning in module (c). Finally, in module (d), shape, texture, and lighting correction modules can be updated accordingly with the instruction from the language brush.
@@ 3 Method
Given a monocular video that captures the human body, we propose an efficient technique to reconstruct the 3D human avatar, which can be further driven by a single-sentence instruction. The overview of our method is shown in Fig. 2, where our approach consists mainly of three fundamental components: 1) we learn a hybrid neural representation to model the human avatar, with a neural feature volume field and neural texture extraction. 2) We introduce a lighting correction network to refine the learned neural texture to cope with lighting differences. 3) In the end, we update the editable neural human representation with language instruction as a brush. With our full pipeline, we can lift the human into a 3D editable neural avatar from a monocular human video alone and edit it with human language descriptions.
@@ 3.1 Reality’s Canvas: Dynamic Human Modeling
$F_{c}:(\mathbf{x}_{c},\mathbf{P},\mathbf{d})\mapsto\mathbf{c},$ (1)
where $\mathbf{x}_{c}$ is a sampled 3D point from the human body in canonical space, $\mathbf{P}$ denotes the human pose for the given time frame, and $\mathbf{d}$ represents the viewing direction. $\mathbf{c}$ is the mapped color in texture space output by the mapping function, $F_{c}$.
@@
where $G_{c}^{F}$ serves as the mapping function to take the 3D input position $\mathbf{x}_{c}$ to obtain the $\mathbf{f}$ to form the neural feature volume.
@@
$G_{c}^{U}:\mathbf{f}\mapsto\mathbf{u},\;\;G_{c}^{V}:\mathbf{f}\mapsto\mathbf{v}$ (3)
@@
$G_{c}^{S}:\mathbf{f}\mapsto\mathbf{s}$ (4)
To model the dynamics in human motion and animate the lifted 3D human avatar, we further introduce the deformation module, $T$, to connect the 3D points $\mathbf{x}$ in the observation space, with the one in the canonical space, $\mathbf{x}_{c}$. In other words, the mapping function in the observation space is connected to the aforementioned mapping function in canonical space via $T$. Thus, we can obtain the neural volume in observation space, $G^{S}$, by warping the canonical neural volume:
$\left(G_{c}^{S}\circ G_{c}^{F}\right)\big{(}T(\mathbf{x},\mathbf{P},\mathbf{d})\big{)},$ (5)
We apply the same deformation field to bend the rays of the mapping function for UV representations. Volume rendering is then applied to the output $\mathbf{U}$, $\mathbf{V}$ and the semantic map $\mathbf{S}$.
@@ 3.2 Disentangled Neural Texture Learning
With the rendered UV map and semantic map, our network can further synthesize high-quality neural textures. We realize this by decomposing appearance generation into diffused color and lighting correction components within a disentangled neural texture learning module. Specifically, we build another neural mapping function, $H_{c}^{T}$, to generate the neural texture field in the texture space with regard to the input UV and semantic space:
@@
Compared to learning a neural radiance field to synthesize the color of the human avatar, learning a neural texture will produce better appearance modeling by avoiding all the floaters above the texture surfaces. However naive neural texture cannot reflect the lighting change affected by occlusion, viewpoint, or human pose alteration. To cope with this issue, we then decompose the neural texture representation into diffused color and lighting correction branches. Firstly, we use MLPs to convert the neural texture representation to the RGB color component, reflecting the diffused color of the surface to represent the human appearance in the bright and evenly distributed lighting condition:
@@
where $\mathbf{a}$, denotes the diffused color component and, $H_{c}^{A}$, represents the mapping function in the diffused color branch.
We further encode lighting correction information to represent the illumination condition influenced by the pose articulation and view direction change:
@@
where $H_{c}^{N}$ is the learned neural mapping function, that maps the neural texture, $\mathbf{t}$, with the guidance of view direction $\mathbf{d}$, and human poses $\mathbf{P}$, to the lighting correction information, $\mathbf{m}$. To this end, the final color for each ray in 2D image plane is calculated as $\mathbf{c}=\mathbf{a}\cdot\mathbf{m}$.
@@ 3.3 Languages’s Brush: Editing Neural Human
Our disentangled neural texture generation allows us to manipulate texture and shape separately. This gives us explicit control over which part of the pipeline is changed for synthesis upon user inputs. Given text instruction for neural human editing, our framework can be customized to update the corresponding module accordingly related to the instruction, enabling more accurate and efficient neural human editing.
@@ 3.4 Training Objectives
where $\lambda_{\text{LPIPS}}$ and $\lambda_{\text{MSE}}$ are the weights for the corresponding losses. The photometric loss will also back-propagate to train the mapping functions, $G_{c}^{U}$, $G_{c}^{V}$, and $G_{c}^{S}$ to predict meaningful UV and semantic maps, which are further regularised by another three losses, $\mathcal{L}_{reg}^{U}$, $\mathcal{L}_{reg}^{V}$ and $\mathcal{L}_{reg}^{S}$ between the predictions and pseudo labels, obtained from a pre-trained Densepose[12] network. In addition to the regularization losses, we add a smoothness term $\mathcal{L}_{\text{smt}}$ to encourage the neighbor 3D points to embed the same semantics by leveraging human parsing priors.
@@
where both $\mathcal{L}_{\text{reg}}^{U}$ and $\mathcal{L}_{\text{reg}}^{V}$ are formulated with MSE loss, $\mathcal{L}_{\text{reg}}^{S}$ uses cross entropy loss, and $\mathcal{L}_{\text{smt}}$ uses $l_{2}$ loss.
To reduce flying pixels and condense the learned neural volume in line with the human shape, we utilize a mask loss to constrain geometry reconstruction with $\mathcal{L}_{\text{mask}}$, similar to [72]. Finally, we obtain the overall training objective in Eq. 11 by weighting all the aforementioned losses to generate the neural human avatar:
@@
where $\lambda_{\text{rec}}$, $\lambda_{\text{reg}}$ and $\lambda_{\text{mask}}$ are the loss weight for $\mathcal{L}_{\text{rec}}$, $\mathcal{L}_{\text{reg}}$ and $\mathcal{L}_{\text{mask}}$, respectively.
In order to train the text editing module, we employ the predicted image as the input, utilizing the ground truth image as the original reference. Editing is performed using InstructPix2Pix [6] in response to the given text prompt. To facilitate this process, the edited image is utilized for computing $\mathcal{L}_{\text{rec}}$, in conjunction with the predicted image.